Cost of Common Equity and WACC

Palencia Paints Corporation has a target capital structure of 40% debt and 60% common equity, with no preferred stock. Its before-tax cost of debt is 10% and its marginal tax rate is 40%. The current stock price is P0 = $23.50. The last dividend was D0 = $3.75, and it is expected to grow at a 7% constant rate. What is its cost of common equity and its WACC? Round your answers to two decimal places. Do not round your intermediate calculations.

  1. rs =  %
  2. WACC =  %
Answer #1

Cost of Common Equity can be calculated using constant growth dividend discount model.

This question requires application of constant growth dividend discount model according to which: Po - Divi T-9 Po = Price of

Div1 = $3.75 * (1 + 7%) = $4.0125

23.50 = 4.0125 r - 0,07

r - 0.07 = 0.1707

Cost of Equity, r = 24.07% (Answer a)

WACC = Weight of debt * Pretax cost of debt * (1 - Tax) + Weight of Equity * Cost of Equity

WACC = 40% * 10% * (1 - 40%) + 60% * 24.07%

WACC = 16.84%(Answer b)